OpenStudy (anonymous):

What is the seventh term in the addition pattern that begins 17, 32, 47, 62? A. 107 B. 117 C. 122 D. 137

OpenStudy (anonymous):

Subtract 32-17=15 32+15=47 47+15=62

OpenStudy (anonymous):

P.s. I'm a big fan of the walking dead also.

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

The answer is a